OpenStudy (anonymous):

After his final exam, a student goes to the top of a building and throws his book into the air. The height h in feet of the book above the ground is given by: h = −16t 2 + 32t + 128. How long is the book in the air?

OpenStudy (anonymous):

it ceases being in the air when it hits the ground. hits the ground when \[h=0\] so set \[-16t^2+32t+128=0\] and solve for t

OpenStudy (anonymous):

i would divide first by -16 to solve \[t^2-2t-8=0\] \[(t+2)(t-4)=0\] \[t=4\]